OpenStudy (anonymous):

FAN AND MEDAL A chef draws cookies randomly from a box containing 6 cookies of the same shape and size. There is 1 chocolate cookie, 3 almond cookies, and 2 butter cookies. He draws 1 cookie and then draws another cookie without replacing the first one. Find the probability of picking 1 almond cookie followed by another almond cookie, and show the equation used.

OpenStudy (anonymous):

I think it is 3/6 * 2/5 = 6/30 =1/5

OpenStudy (anonymous):

I got the same thing so I think it is correct see.. 3/6 * 2/5 = 6/30 =1/5

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Welcome